# The Software Diversity Card: A Framework for Reporting Diversity in Software Projects

### v1

@@ 3.2 Study results
4 Dimension of the diversity card
This section provides a comprehensive overview of the various components of the Software Diversity Card, presenting a structured profile of the different groups identified in Section 2. The card is organized into three main parts: the first part profiles the collection of teams involved in the software project, highlighting not only the developers but also non-coding contributors and testers; the second part outlines the usage context, detailing the intended target communities and any adaptations made for specific social groups; and finally, the last part describes the governance processes along with the relevant government bodies (such as funders and shareholders) that influence the project. Figure 1 provides an overview of the card’s structure.
@@ 4.1 Participants in software projects
This part aims to profile the different types of participants behind developing and maintaining a software project. For each of these teams, and inspired by recent works in software diversity and inclusion [1, 2], the goal is to document a minimum set of demographic attributes such as age, perceived gender, ethnicity, location, and disabilities; and cultural characteristics of the team, such as the spoken language, the socio-economic status, the job title, the experience, and education level, among others. The clearest team behind a system is the development team. Several studies have studied how diverse teams are beneficial for the outcome of software systems, and this information would be valuable for end-users to trust in the software (e.g., the case of the period tracker app), or for public regulators to ensure the app is developed in a certain region
However, there are other types of teams involved in the development of software products that also play a pivotal role. Non-coding contributors have been recently noticed as relevant in the evolution and maintenance of open-source software projects [15]. For example, roles like the issue reporters are influential in the evolution of software projects. However, social groups less familiar with reporting issues may see their needs under-represented during the evolution of software systems [37]. To this end, gathering diverse information about reporters could help companies avoid this kind of bias.
@@ 4.2 The usage context of the software
To properly understand why a specific composition of actors is relevant, we need information about the social context where a system is intended to be used. Our proposal allows reporting the particular use cases the software is designed for and the social and cultural constraints that are applicable to these use cases. Following the example of the period tracker app, this app could be designed for young women living in Catalonia, a region of Spain with a specific co-official language (Catalan). This social context may generate the need to adapt the app’s user interface in Catalan and test the app on Catalan speakers to ensure it fits their needs.
@@ 5 Modeling diversity in software projects
In this section, we first present the abstract syntax of the DSL, which formalizes the concepts outlined in Section 4. Then, we introduce two different concrete syntaxes, one tailored for users with coding skills, implemented as a language plugin for Visual Studio Code, and a second one oriented to non-coders based on a web-based tool that allows non-technical people to create cards easily.
@@ 5.1.1 Participants
Figure 2: Metamodel for collecting data about generic entities and individuals.
Individuals represent both team participants and other individual stakeholders in software projects. Individuals report their spokenLanguages (including their respective proficiency levels, based on the Common European Framework of Reference for Languages [^16] https://www.coe.int/en/web/common-european-framework-reference-languages), ethnicity, gender, age, socioEconomicStatus (like the Kuppuswamy socioeconomic status class classification [54] and similar), skillLevel and tenure, among other information. On the other hand, Groups include homologous data, but in aggregated format, i.e., a project reports an aggregated profile of the development team instead of the profile of each individual. In that sense, Groups represent the essential piece of our model, and Teams, Organizations, and TargetCommunities are the specific types of Groups that compose the card.
In Figure 3, we can see an excerpt of the team hierarchies and its participants. As per their objectives and responsibilities, Teams are classified as: TesterTeams, PublicReporterTeams, NonCodingContributorTeams, or DevelopmentTeams; each of them containing specific information. Teams can be internal for the software project, or can be hired via a labour force. For instance, the TesterTeams can be hired via outsourcing services. In that case, Teams are classified as LabourForces. These include further details regarding contractual aspects, such as salary and labourRights (primarily according to the Country where the team is based in), and the company the team belongs to.
@@
Teams are integrated by Participants, who belong to a Team within a concrete date range, and exercise a particular role (e.g., reporter, developer or reviewer, among others). Those Participants who are affiliated to the organization responsible for the development may inform their participantId (i.e., their internal business identifier; e.g., to enable them to access the organization’s owned systems).
@@ 5.1.2 Usage context
As mentioned in previous sections, software systems must consider its usage context, captured in Figure 4. In that sense, the SocialContext reflect the environment of social interactions, spokenLanguages, and culturalTraits where a software system is going to be exploited, in order to effectively address the users’ needs and nuances. The features of a system are built to respond to the different intended UseCases and their respective TargetCommunities or end users. For particular TargetCommunities who have specific needs or alternative, complementary requirements, it may be necessary to make additional Adaptations delivered in a release.
@@ 5.2.2 A form-based web tool for creating software diversity cards
6 Examples
@@
In this section, we describe two mature open-source projects involving different groups of people using the proposed software diversity card. The two projects are Decidim [56], a participatory software used by over 400 administrations worldwide to allow citizens to participate in public decisions, and Besser [57], a low-code platform designed to enable non-coding profiles to develop software applications in various domains which one of the authors of this work is directly involved. Despite their distinct characteristics, such as target users, both projects are open-source, publicly funded, and part of a user community.
@@
As a first step in describing the projects, we reviewed the online documentation of each project and attempted to complete the cards as much as possible. Then, we interviewed the project’s development and product teams, presenting our proposal and asking about the card’s dimensions. This section summarizes the key points of each project’s card and discusses the challenges we and the software projects encountered in gathering the required card dimensions. The complete examples of the cards are available online [^22] The card’s examples are in the /examples folder of the tools repository: https://github.com/SOM-Research/SoftwareDiversityCard.
@@ 6.1 Decidim: an open-source participatory democracy platform
Decidim is an open-source platform that fosters participatory democracy and involves different groups participants. The platform’s development team comprises maintainers, developers, and external contributor teams focused on specific features. Despite the extensive project documentation and clear commitment to transparency, we have found that the developers’ profiles were not publicly disclosed. During the interview with Decidim, we learned that Decidim is composed of internal and external developer teams. In that sense, they pointed out that the external teams change over time, making it difficult for them to publish a long-standing team’s profile information. The description of the current state of the Decidim development team can be seen in the online example of the Decidim Cards.
Decidim has a community of non-coding contributors who report bugs, propose features, and translate the software via its own online social platform, Metadecidim [^23] Metadecidim home page: https://meta.decidim.org. However, following the approach of many open-source communities, they adopted the “anonymity first” principle, where any user contribution only needs to be associated with a user name. While this approach is intended to avoid bias and promote equal participation, it is hard for them to profile the community of non-coder contributors. In that sense, we have been able to profile the size, number of contributions, and its expertise in the reporting platforms of both teams, that can be seen in the complete example published online.
Regarding testers and public reporters, they did several usability tests and public interviews with citizens in several participatory processes they drive. However, none of these studies have been made public or reported in the extensive project’s documentation. They noted that this is a path of improvement for the project, as disclosing this information may help, for instance, in the Decidim’s adoption by other institutions. However, they pointed to some limitations regarding this issue, such as the lack of UI/UX experts to build systematic tests and standard guidelines to disclose this information in open-source projects.
Listing 3: Excerpt of the usage context part of the Decidim’s Card
@@
They have also conducted several accessibility audits to adapt the software for people with disabilities. However, while various governments have conducted these audits, the results remain unpublished. During the interview, they commented that one of the reasons this information is not published is the lack of resources and clear guidance on how to publish these results, which are sometimes hard for final users to understand as one of the reasons. However, they agree that pointing to the results of audits would be useful for other governments willing to adopt Decidim in their participatory processes.
@@ 6.2 Besser: a low-code platform for smart software development
Besser is a low-code platform that allows non-coding developers to develop domain-specific applications. It is maintained and developed by a community of participants; for instance, the development team is composed of 15 developers from 9 nationalities, where French, English, and Spanish are the running languages, and the educational level is balanced between undergraduates and doctoral degree holders in computer science. While this information is partially public on the project website, one of the authors of this work is directly involved in developing this software, so we had direct access to the development team and the information required. Still, we needed to gather the geographic distribution, the running languages, and the education level from the interviews with the project managers. In Listing 4 we can see the different attributes captured in the Besser diversity card.
@@
Listing 4: Excerpt of the development team part of the Besser’s Card
@@
Regarding the usage context, the adaptations are one of the key elements of Besser, as it is a low-code platform intended to help users develop domain-specific software. In that sense, the software presents a set of domain adaptations identifying a target user for each domain. For instance, they created an extension of Besser to allow public servants and researchers to create visualization panels for climate data; on the other hand, they adapted Besser for scholars, focusing on teachers and students in computer science, and an ongoing project involves adapting Besser to telecommunications engineers to model the deployment of 6G networks. Finally, regarding governance, Besser is a public-funded project funded by the FNR Pearl grant led by the Luxembourg Institute of Science and Technology. The principal investigator is the grant recipient and acts as a project director (unlike Decidim, which is governed by a board of directors). Besser’s governance profile can be seen in Besser’s online published card.
7 Conclusions
In this work, we reviewed the increasing concerns within the research field regarding diversity in software projects. These concerns pertain not only to the development team but also to a broader definition of participants, which includes non-coding roles, testers, and intended end-users. Despite the growing awareness of diversity in the research community, our analysis of 1,000 popular open-source software (OSS) projects revealed a significant lack of reported diversity information.
To address this issue, we proposed the Software Diversity Card as an integrated approach to report diversity-related aspects within this broader definition of participation. Our proposal aims to serve as a communication tool for software owners, users, and public regulators. To facilitate its adoption, we have also formalized a domain-specific language (DSL) along with a toolkit to assist in creating the card. This formalization provides a common language that enables stakeholders involved in a software project to understand, communicate, and discuss diversity-related aspects effectively.
Our proposal represents an initial effort to acknowledge diversity as a crucial element in software projects. However, there are still challenges to address and future directions to explore in this area. In the remainder of this section, we will outline some of the challenges we faced during the description of Decidim and Besser and propose future research avenues focused on reporting diversity-related aspects.
@@ 7.1 Challenges
Anonymity: One of the main limitations we faced while creating the Decidim and Besser card was anonymity. While crucial for reducing biases, anonymity in open-source communities poses significant challenges in profiling contributors. In the case of Decidim, the ”anonymity first” principle helps prevent discrimination against specific groups but complicates the process of analyzing the non-coding contributor’s teams, such as the translators or the reporters of the software. This approach limits the ability to conduct a post-analysis of the community’s composition, making it harder to effectively understand and support the broader user base.
Sensitive information: Related to anonymity, another clear challenge is the presence, in the software diversity card, of sensitive information about a project’s developers, contributors, and users, including details such as racial or ethnic background, age, gender, and sexual orientation. While this information is generally provided in aggregate form, in smaller software projects, even aggregate data can potentially be used to infer details about individuals. This raises several privacy concerns for the software diversity card. For instance, surveys that collect sensitive information from individuals must include informed consent, clearly stating what data is being collected and how it will be used. On the other hand, if the population size or characteristics of the data allow for individual attribution of traits, such information should not be included in the software diversity card.
@@
Authenticity: The software diversity card is a self-assessment report and, as such, it assumes goodwill on the part of the authors in terms of its authenticity. Nevertheless, the creators of a diversity card may have strong incentives to misrepresent the level of diversity in a software project, e.g., to attract more talent, to access funding from public organizations or specific communities, to gain publicity or public support, or to strengthen its participation in competitive calls. Thus, it is important to emphasize that the purpose of the software diversity card is to improve community diversity practices in software development. Such a goal can only be achieved if the information presented in diversity cards is accurate.
@@
Evolving nature of projects: One of the challenges during the creation of Decidim’s card has been the evolving nature of the development team. The team comprises internal and external teams, which could easily change shortly. They pointed out that this situation makes disclosing the team’s long-standing profile information difficult. The same problem could happen when new tests are done with users, or the evolution of the reporting community. So, the inherent evolving nature of the project forces us to frequently reanalyze the card, which can be helpful to determine whether the project is improving or worsening in terms of diversity and inclusion.
7.2 Future work

### v2

@@ 3.3 Threats to validity of the study
4 Dimensions of the diversity card
This section introduces the structure of the Software Diversity Card, which we organized into three main parts; Participants, Usage Context, and Governance. In Figure 1 we can see an overview of the card structure.
The proposed separation is proposed as a way of distinguishing between endogenous and exogenous determinants of software projects within a socio-technical system [45]. Participants correspond to endogenous agents whose activities, decisions, and interactions are generated within the boundaries of the project and directly affect its evolution, while the Usage Context (end users and broader social environment) corresponds to exogenous factors that constrain or shape project outcomes. Governance serves as a mediating layer, operating as a boundary-setting mechanism that connects internal agency with external conditions [67]. This tripartite distinction aligns with long-standing perspectives in socio-technical systems research [68, 69, 45]), which emphasize that successful systems design requires balancing the technical core (the participants’ actions), the social context (usage context), and governance.
@@ 4.1 Participants in software projects
This part aims to profile the different types of participants behind developing and maintaining a software project. For each of these teams, and inspired by recent works in software diversity and inclusion [1, 2], the goal is to document a minimum set of demographic attributes such as age, perceived gender, ethnicity, location, and disabilities; and cultural characteristics of the team, such as the spoken language, the socio-economic status, the job title, the experience, and education level, among others. The most clearly identifiable team behind a system is the development team. Several studies have studied how diverse teams are beneficial for the outcome of software systems, and this information would be valuable for end-users to trust in the software (e.g., the case of the period tracker app), or for public regulators to ensure the app is developed in a certain region
However, there are other types of teams involved in the development of software products that also play a pivotal role. Non-coding contributors have largely been noticed as relevant in the evolution and maintenance of open-source software projects [17, 28]. For example, roles like the issue reporters are influential in the evolution of software projects. However, social groups less familiar with reporting issues may see their needs under-represented during the evolution of software systems [49]. To this end, gathering diverse information about reporters could help companies avoid this kind of bias.
@@ 4.2 The usage context of the software
To properly understand why a specific composition of actors is relevant, we need information about the social context where a system is intended to be used. Our proposal allows reporting the particular use cases the software is designed for and the social and cultural constraints that are applicable to these use cases. Following the example of the period tracker app, this app could be designed for young women living in Catalonia, a region of Spain with a specific co-official language (Catalan). This social context may generate the need to adapt the app’s user interface to Catalan and test the app on Catalan speakers to ensure it fits their needs.
@@ 5 Modeling diversity in software projects
In this section, we first present the abstract syntax of the DSL, which formalizes the concepts outlined in Section 4
Then, we introduce two different concrete syntaxes, one tailored for users with coding skills, implemented as a language plugin for Visual Studio Code, and a second one oriented to non-coders based on a web-based tool that allows non-technical people to create cards easily.
@@ 5.1.1 Participants
Figure 2: Metamodel for collecting data about individuals and groups.
Individuals are a generic representation of the common data to be collected for both team participants and other individual stakeholders in software projects. Individuals report their spokenLanguages (including their respective proficiency levels, based on the Common European Framework of Reference for Languages [^13] https://www.coe.int/en/web/common-european-framework-reference-languages), ethnicity, gender, age, socioEconomicStatus (like the Kuppuswamy socioeconomic status class classification [73] and similar), skillLevel and tenure, among other information. On the other hand, Groups include homologous data, but in aggregated format, in order to uniquely specify the generic data that is collected for representing groups of individuals. For instance, a project may report an aggregated profile of the development team instead of the profile of each of their individual members. In that sense, Groups represent the essential piece of our model, and Teams, Organizations, and TargetCommunities are the specific types of Groups that compose the card.
In Figure 3, we can see an excerpt of the team hierarchies and its participants. Teams are classified into two orthogonal specialization sets. The first set specializes a Team according to its objectives and responsibilities, capturing its functional role within the system. Teams are therefore classified as either TesterTeams, PublicReporterTeams, NonCodingContributorTeams, or DevelopmentTeams; each of them contributing with specific information. The second specialization set classifies Teams based on organizational sourcing, distinguishing between internal and external LabourForces, for which further details regarding contractual aspects such as salary and labourRights (primarily according to the Country where the team is based in), and the company the team belongs to is informed. These two specialization sets are independent, in order to comprehend diverse combinatorial scenarios. For instance, a TesterTeam could be outsourced as a LabourForce, whereas a DevelopmentTeam could be internal for the development of a software system.
@@
In addition, Participants are a specialization of Individuals, so as to distinguish the concrete actors that are associated with Teams. Participants belong to a Team within a concrete date range, and exercise a particular role (e.g., reporter, developer or reviewer, among others) in such period. Those Participants who are affiliated to the organization responsible for the development may inform their participantId (i.e., their internal business identifier; e.g., to enable them to access the organization’s owned systems) and are represented as InternalParticipants.
@@ 5.1.2 Usage context
As mentioned in previous sections, software systems must consider its usage context, captured in Figure 4. In that sense, the SocialContext reflects the environment of social interactions, spokenLanguages, and culturalTraits where a software system is going to be exploited, in order to effectively address the users’ needs and nuances. The features of a system are built to respond to the different intended UseCases and their respective TargetCommunities or end users. For particular TargetCommunities who have specific needs or alternative, complementary requirements, it may be necessary to make additional Adaptations delivered in a release.
@@ 5.2.2 A form-based web tool for creating software diversity cards
5.3 DSL evaluation
To evaluate the usefulness, completeness, and usability of the proposed DSL and its supporting tools, we conducted an empirical validation with experienced software developers. The experiment comprised a reading and a writing exercise. Initially, participants used the VSCode language plugin to read a complete diversity card and answered questions to measure the readability of the DSL. Subsequently, using the form-based web tool, they completed a partially filled-out diversity card, accompanied by a document containing project diversity information in natural language. This part assessed the usability of the tool and the identification of complex concepts, such as target communities and user reporting platforms. Post-exercises, participants took part in a semi-structured interview, providing quantitative feedback via Likert-scale items across criteria such as (1) usefulness (ability to capture diversity information), (2) completeness (adequacy of the DSL constructs), and (3) usability (ease of use and adoption potential); and qualitative insights through open questions within each category.
The experiments took place in August and September 2025 in Spain and Luxembourg, with 13 English-fluent participants composed mainly of research software developers from three research groups, with a median of four years of experience. The complete questionnaire and detailed study results are available in an online repository [^19] The documentation can be found in the docs folder of the tools repository: https://github.com/SOM-Research/SoftwareDiversityCard.
The reading exercise results show that 93.5% of responses were correct, indicating a high level of DSL readability. Most participants successfully identified key concepts such as target communities and reporting platforms. Incorrect answers were primarily provided by non-developers, which suggests that limited experience with VSCode may have introduced a bias. Divergent interpretations emerged around governance: while some participants recognized funders (e.g., the Barcelona Council) as relevant actors, others referred only to the project’s board.
Figure 8: Results of the questionnaire grouped by the assessment criteria. Scores range from 1 (strongly disagree) to 5 (strongly agree). Usefulness comprises Q1, Q2, and Q3; Completeness comprises Q4, Q5, Q6; and Usability comprises Q7, Q8, and Q9 of the questionnaire.
The results of the writing exercise and post-exercise interviews are summarized in Figure 8. The Likert-scale questions, grouped by evaluation criteria, reflect a positive assessment of the DSL and its supporting tools, with mean scores above 4 (on a 1–5 scale) for usefulness, completeness, and usability. In contrast, responses to the open-ended questions yielded several areas for improvement. Participants frequently mentioned compatibility issues with browsers and mobile devices, underscoring the need to adapt the tool for diverse platforms. They also recommended adding intersectionality options (e.g., combining gender and race) to balance privacy with informativeness. Additional suggestions included enriching the tool with graphical elements, such as maps, to support visualization of geographic distribution, and linking user profiles to specific roles in software projects (e.g., managers or developers). The latter could be realized through integration with business process management tools, as discussed in Section 8.
6 Use cases
@@
In this section, we describe two mature open-source projects involving different groups of people using the proposed software diversity card. The two projects are Decidim [75], a participatory software used by over 400 administrations worldwide to allow citizens to participate in public decisions, and Besser [76], a low-code platform designed to enable non-coding profiles to develop software applications in various domains where one of the authors of this work is directly involved. Despite their distinct characteristics, such as target users, both projects are open-source, publicly funded, and part of a user community.
@@
As a first step in describing the projects, we reviewed the online documentation of each project and attempted to complete the cards as much as possible. Then, we interviewed the project’s development and product teams, presenting our proposal and asking about the card’s dimensions. This section summarizes the key points of each project’s card and discusses the challenges we and the software projects encountered in gathering the required card dimensions. The complete examples of the cards are available online [^20] The card’s examples are in the /examples folder of the tools repository https://github.com/SOM-Research/SoftwareDiversityCard, and Tables 2 and 3 summarize the overall completeness of the card for each use case.
@@ 6.1 Decidim: an open-source participatory democracy platform
Decidim is an open-source platform that fosters participatory democracy and involves different groups participants. The platform’s development team comprises maintainers, developers, and external contributor teams focused on specific features. Despite the extensive project documentation and clear commitment to transparency, we have found that the developers’ profiles were not publicly disclosed. During the interview with Decidim, we learned that Decidim is composed of internal and external developer teams. In that sense, they pointed out that the external teams change over time, making it difficult for them to publish a long-standing team’s profile information. This evolving nature of the development team represents a challenge for the proposed card, in that specific case, we decided to capture the current state of the team, which can be seen in the complete example published online.
Decidim has a community of non-coding contributors who report bugs, propose features, and translate the software via its own online social platform, Metadecidim [^21] Metadecidim home page: https://meta.decidim.org. However, following the approach of many open-source communities, they adopted the “anonymity first” principle, where any user contribution only needs to be associated with a username. While this approach aims to avoid bias and promote equal participation, it is challenging for them to profile the community of non-coder contributors accurately, and it represents an open challenge for the Software Diversity Card. In that sense, we have been able to profile the size, number of contributions, and its expertise in the reporting platforms of both teams, that can be seen in the complete example published online.
Regarding testers and public reporters, they did several usability tests and public interviews with citizens in several participatory processes they drive. However, none of these studies have been made public or reported in the extensive project’s documentation. They noted that this is a path of improvement for the project, as disclosing this information may help, for instance, in the adoption of Decidim by other institutions. However, they pointed to some limitations regarding this issue, such as the lack of UI/UX experts to build systematic tests and standard guidelines to disclose this information in open-source projects. This example illustrates one of the benefits of standardization efforts in this area, where the lack of guidance hinders the disclosure of diversity-related information.
Table 2: Overview of the completeness of Decidim’s card.
Dimensions Team type Profile completeness
Participants Developer Team High as a summary of internal and external teams
Non-coding contributors Limited due to OSS community (anonymity-first)
Testers High as extracted from public interviews
Reporters Limited due to the anonymity-first approach
Usage Context Target Communities Limited to the project’s use cases
Governance Boards High as is public information
Bodies High as they are public institutions
Listing 3: Excerpt of the usage context part of the Decidim’s card.
@@
They have also conducted several accessibility audits to adapt the software for people with disabilities. However, while various governments have conducted these audits, the results remain unpublished. During the interview, they commented that one of the reasons this information is not published is the lack of resources and clear guidance on how to publish these results, which are sometimes hard for final users to understand as one of the reasons. However, they agree that pointing to the results of audits would be useful for other governments willing to adopt Decidim in their participatory processes. Similarly to the usability test, this shows a benefit of the standardization efforts like our proposal, which could guide on how to disclose accessibility information.
@@ 6.2 Besser: A low-code platform for smart software development
Besser is a low-code platform that allows non-coding developers to develop domain-specific applications. It is maintained and developed by a community of participants; for instance, the development team is composed of 15 developers from 9 nationalities, where French, English, and Spanish are the running languages, and the educational level is balanced between undergraduates and doctoral degree holders in computer science. Although this information is partially public on the project website, the interviews with the team have provided us with direct access to the development team and the necessary information. Still, we needed to gather information on the geographic distribution, the languages used, and the education levels from the interviews with the project managers. Listing 4 shows the different attributes captured in the card, and Table 3 shows an overview of the card’s completeness.
@@
Table 3: Overview of the completeness of Besser’s card.
Dimensions Team type Profile completeness
Participants Developer Team High, reported directly by development team
Non-coding contributors Limited due to anonymity-first approach of OSS
Testers High, tests lead by software owners
Reporters Limited due to anonymity-first approach of OSS
Usage Context Target Communities Limited to the project’s use cases
Governance Boards High, principal investigator
Bodies High, public institutions
Listing 4: Excerpt of the development team part of the Besser’s card.
@@
Regarding the usage context, the adaptations are one of the key elements of Besser, as it is a low-code platform intended to help users develop domain-specific software. In that sense, the software presents a set of domain adaptations identifying a target user for each domain. For instance, they created an extension of Besser to allow public servants and researchers to create visualization panels for climate data; on the other hand, they adapted Besser for scholars, focusing on teachers and students in computer science, and an ongoing project involves adapting Besser to telecommunications engineers to model the deployment of 6G networks. Regarding governance, Besser is a public-funded project funded by the FNR Pearl grant led by the Luxembourg Institute of Science and Technology. The principal investigator is the grant recipient and acts as a project director (unlike Decidim, which is governed by a board of directors).
7 Discussion
In this section, we discuss the benefits of the proposed Software Diversity Card for both practitioners and the research community, then we discuss the challenges that emerged during its application, and finally, we present the threats to the validity of our study.
For practitioners, the card represents an initial effort to recognize diversity as a critical dimension of software projects. In the applied use cases, the card enabled us to collect information that was not readily available in public documentation, despite its potential to enhance transparency and project value. Additionally, the card provided guidance on which diversity-related information is important to share and offered a structured format for doing so. For example, in the Besser and Decidim projects, usability and accessibility tests were routinely conducted for internal purposes but rarely published, even though sharing these results could encourage adoption by institutions and public administrations, while also promoting transparency and accountability. Overall, the card serves as a practical tool for making diversity considerations more visible in software projects by practitioners. By highlighting key information and offering a structured approach to disclosure, it has the potential to enhance diversity-sharing practices and serve as a communication tool between practitioners, users, and public institutions.
For the research community, this initiative aims to stimulate greater attention to diversity in software development. Analogous to reproducibility badges, the Software Diversity Card could serve as a formal mechanism to encourage the inclusion of diversity considerations in submission guidelines and review procedures for scientific venues, particularly those publishing software-related work. By providing a structured way to report and reflect on diversity aspects, the card can help raise awareness among researchers and reviewers, fostering a culture that values and actively addresses diversity. We also recognize that diversity, like any social phenomenon, is dynamic and evolves alongside the societies in which software is developed and deployed. The card should therefore be viewed as an initial step rather than a definitive solution. Its adoption and refinement over time can support ongoing reflection, dialogue, and iterative improvements in how diversity is understood, measured, and promoted within the software development and research communities.
@@ 7.1 Challenges
Several challenges have emerged during the application of the card. One of the main challenges encountered when creating the Decidim and Besser cards was the anonymity of the contributors. While anonymity is essential for reducing biases and protecting participants, it poses significant challenges for profiling contributors in open-source communities. For example, in Decidim, the “anonymity first” principle helps prevent discrimination against specific groups but complicates the analysis of non-coding contributor teams, such as translators or software reporters. This limitation restricts the ability to perform post-hoc analyses of community composition, making it more difficult to understand and support the broader user base effectively.
Closely related to anonymity is the challenge of handling sensitive information in the Software Diversity Card, including data on developers, contributors, and users, such as race or ethnicity, age, gender, and sexual orientation. While such information is typically aggregated, in smaller projects, even aggregate data may allow for individual identification, raising privacy concerns. Surveys and data collection processes must therefore include informed consent, clearly communicating what data is collected and how it will be used. Moreover, when population size or characteristics permit the inference of individual traits, this information should not be included in the card. These issues underscore the importance of aligning the Software Diversity Card with current privacy regulations, such as GDPR. Ensuring compliance while capturing meaningful diversity-related insights remains an open challenge and represents a key direction for future research on the practical application and governance of the card.
@@
The software diversity card is a self-assessment report and, as such, it assumes goodwill on the part of the authors in terms of its authenticity. Nevertheless, the creators of a diversity card may have strong incentives to misrepresent the level of diversity in a software project, e.g., to attract more talent, to access funding from public organizations or specific communities, to gain publicity or public support, or to strengthen its participation in competitive calls. One potential mechanism to alleviate this challenge is the inclusion of a third-party system to store and share card information, acting as a trust partner for external users. It is important to emphasize that the purpose of the software diversity card is to improve community diversity practices in software development. Such a goal can only be achieved if the information presented in diversity cards is accurate.
@@
One of the challenges during the creation of Decidim’s card has been the evolving nature of the development team. The team comprises internal and external teams, which could easily change shortly. They pointed out that this situation makes disclosing the team’s long-standing profile information difficult. The same problem could happen when new tests are done with users or the evolution of the reporting community. So, the inherent evolving nature of the project forces us to frequently reanalyze the card, which can be helpful to determine whether the project is improving or worsening in terms of diversity and inclusion.
Practitioners may encounter several challenges when adopting the Software Diversity Card. A primary concern is the additional effort required, which can be a significant barrier for small- and medium-sized projects, while larger projects may be better equipped to manage this complexity. This creates a potential risk of restricting competition, as smaller projects could be excluded from public tenders due to the administrative burden, which is both time-consuming and resource-intensive. To alleviate this issue, we developed a toolkit to support the use and creation of the card; however, addressing this challenge fully will require a mature toolkit ecosystem that evolves alongside the needs of the software development community. An interesting direction for tool development is the integration of the Software Diversity Card with existing documentation assets, such as README files, codes of conduct, and governance documents. This integration could be achieved either by automatically generating or updating documentation from a valid card, or by populating the card by parsing existing documentation—leveraging, for example, the promising capabilities of Large Language Models in feature extraction from technical documents [77]. Another important challenge is the lack of incentives for adopting diversity practices. Beyond the intrinsic benefits of promoting diversity, external mechanisms—such as reproducible badges awarded by public regulators, similar to those used in scientific publications—could serve as tangible incentives for organizations to implement and report on diversity measures.
7.2 Threats to validity
A key limitation of this study is its focus on open-source software projects and their public documentation, which may restrict the generalizability of our findings to other contexts. While open-source communities prioritize transparency, they might have additional documentation not considered here. In contrast, commercial software platforms often limit access to documentation, which could skew the practices seen in open-source projects. Future research should expand the evaluation to include a wider range of software ecosystems and additional documentation sources.
A second threat arises from our use of large language models (LLMs), specifically GPT-4o-mini, for textual analysis and extraction of information. While LLMs have demonstrated strong performance in analyzing and summarizing text, they are prone to errors, omissions, and hallucinations, which could affect the validity of the extracted data. To mitigate this risk, we performed manual validation of the model outputs and cross-checked key findings with the project documentation and interviews. Nonetheless, the potential for subtle inaccuracies remains a consideration when interpreting the results.
Finally, the geographical context of the case studies introduces another potential threat to validity. Both the Besser and Decidim projects are based in Europe, and diversity-related aspects are highly sensitive to regional cultural and social norms. As such, our findings may not generalize to projects situated in other regions with different social, cultural, or regulatory contexts. Addressing this limitation will require future studies to examine software projects in more diverse geographical settings, which would provide a more comprehensive understanding of how diversity manifests across global software development practices.
8 Conclusions and future work
In this work, we examined the growing concerns in the research community regarding diversity in software projects, extending the focus beyond development teams to include non-coding roles, testers, and end users. Despite increasing awareness, our analysis of 1,000 popular open-source projects revealed a substantial lack of reported diversity information. To address this gap, we proposed the Software Diversity Card as an integrated approach for reporting diversity-related aspects across this broader definition of participation. To facilitate adoption, we also formalized a domain-specific language (DSL) and developed a supporting toolkit, providing a common framework for stakeholders to understand, communicate, and discuss diversity considerations effectively.
We validated the card through two real-world case studies, analyzing public documentation and conducting interviews with development teams. The findings illustrate both the benefits and challenges of using the card in practice. Among the benefits, the card enabled the collection of valuable information not publicly documented, such as internal usability test results, which could enhance project transparency and value. The challenges included issues related to anonymity and sensitivity, which limited the collection of nuanced participant data, the dynamic nature of development teams, which complicates capturing their full composition and evolution, and adoption barriers, particularly the additional effort required for small- and medium-sized projects. These results highlight the potential of the Software Diversity Card as a tool for promoting awareness and reporting of diversity, while also indicating areas for further refinement and support to ensure practical applicability across diverse software projects.